OpenStudy (anonymous):

What is the slope of the line that contains the points (0, 3) and (8, -3)? A. -4/3 B. -3/4 C. 4/3 D. 3/4

jimthompson5910 (jim_thompson5910):

Are you familiar with the slope formula? \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

Kinda, I haven't done slopes for a while..

OpenStudy (triciaal):

the slope is the ratio of change in y with change in x as indicated by the formula above

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

Im kind of confused on it,.

jimthompson5910 (jim_thompson5910):

what are the y coordinates of the points (0, 3) and (8, -3)

OpenStudy (anonymous):

3 and -3?

jimthompson5910 (jim_thompson5910):

subtract the two y coords: y2 - y1 = (-3) - (3) = -6 ------------------------------------------------------- then you subtract the x coords in the same order x2 - x1 = (8) - (0) = 8 ------------------------------------------------------- Then you divide the y result by the x result: -6/8 = -3/4 notice how it's y over x (not x over y)

OpenStudy (anonymous):

Yes. So then it would be B.

jimthompson5910 (jim_thompson5910):

correct